OpenStudy (anonymous):

how can i find the coordinates of the midpoint of a segment having the endpoints: Q(1,-3),R(11,5)?

OpenStudy (amistre64):

Q+R ---- 2

OpenStudy (amistre64):

1,-3 11, 5 ----- (12, 2)/2 = (6,1)

OpenStudy (amistre64):

is that right?

OpenStudy (amistre64):

nah, its subtracted and /2 i think would be better

OpenStudy (amistre64):

1+5 = 6 -3+4 = 1 .. nope, that first bit was right lol

OpenStudy (anonymous):

so, everytime they ask me for the coordinates of the midpoint, i just add them and divide by 2?